OpenStudy (anonymous):

The weight of 10 identical samples of a substance is 0.1 pounds What is the weight of 10,000 samples?

OpenStudy (anonymous):

10/.1 = 10,000/x (cross multiply) (10)(x) = (.1)(10,000) 10x = 1000 x = 1000/10 x = 100 The weight of 10,000 samples is 100 lbs.